### v1

# Resource-Theoretic Hierarchy of Contextuality for General Probabilistic Theories

Abstract
In this work we present a hierarchy of generalized contextuality. It refines the traditional binary distinction between contextual and noncontextual theories, and facilitates their comparison based on how contextual they are. Our approach focuses on the contextuality of prepare-and-measure scenarios, described by general probabilistic theories (GPTs). To motivate the hierarchy, we define it as the resource ordering of a novel resource theory of GPT-contextuality. The building blocks of its free operations are classical systems and GPT-embeddings. The latter are simulations of one GPT by another, which preserve the operational equivalences and thus cannot generate contextuality. Noncontextual theories can be recovered as least elements in the hierarchy. We then define a new contextuality monotone, called classical excess, given by the minimal error of embedding a GPT within an infinite classical system. In addition, we show that the optimal success probability in the parity oblivious multiplexing game also defines a monotone in our resource theory. We end with a discussion of a potential interpretation of the non-free operations of the resource theory of GPT-contextuality as expressing a kind of information erasure.
1 Introduction
Generalized (non)contextuality.
A crucial research question in the foundations of quantum theory is to identify those features of quantum theory that constitute a true departure from the classical worldview. Addressing this question requires one to first establish a good notion of classicality, which adequately captures the classical worldview. We believe that a good notion of classicality should satisfy the following desiderata (see also [1]): (1) it endorses a principle that defines a clear boundary between aspects that pose interpretational issues and those that do not, (2) it is of broad range of applicability, (3) it is empirically testable, and (4) its violation constitutes a resource for practical applications, in particular in quantum information processing. [^1] This last desideratum is motivated by the belief that identifying the true nonclassicality of quantum theory will ultimately provide the answer to the question about the origin of the alleged quantum computational speed-up. Motivated by these desiderata, a leading notion of classicality is generalized noncontextuality [2]. [^2] In what follows we will often omit “generalized” and just talk of “(non)contextuality”. A noncontextual theory is one that is compatible with a classical realist explanation of its operational predictions —  it admits of a noncontextual ontological model. In such a model, any two experimental procedures that the theory predicts to be operationally indistinguishable also have the same ontological representation (see Section 2 for a precise definition ).
Noncontextuality satisfies (1), in that it is an instance of a methodological principle inspired by Leibniz’s principle of the identity of indiscernibles [3] (also formulated as a no fine-tuning principle [4]). A violation of such principle would indeed entail an interpretational problem, as it would attribute a conspiratorial connotation to the realist explanation of the theory: why should experimental procedures predicted by the theory to be indistinguishable in principle be represented by different distributions in the ontological model? Noncontextuality also satisfies (2), as it applies to a wide range of scenarios including prepare-and-measure experiments of a single system, unlike Bell’s local causality — another leading notion of classicality. Moreover, in situations where these are applicable, it coincides with notions of classicality such as non-negativity of quasi-probability representations [5, 6] and Bell’s local causality [7]. It satisfies (3), as witnessed by the experiments performed to test quantum violations of generalized noncontextuality [8, 9]. Finally, desideratum (4) has been argued by the numerous works showing that contextuality is a resource for information processing tasks [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21].
Hierarchy of contextuality.
In relation to the last point, contextuality can be used to witness and characterize the advantage of using quantum physical systems in practical applications. In such contexts, it is important to know not just whether a theory is contextual, but also to quantify how contextual it is. Our article aims to address this by introducing a hierarchy of contextuality, in which not all contextual phenomena are equivalent. In particular, it allows us to make more fine-grained distinctions and to quantify the amount of contextuality present in a theory.
To motivate the hierarchy we propose, we take inspiration from the framework of resource theories [22, 23, 24]. There, a resource object $r$ (such as a quantum state) is deemed more valuable than a resource object $s$, denoted $s\preceq r$, if there is a free operation that transforms $r$ to $s$. The preorder relation [^3] A preorder is a relation that is both reflexive and transitive. It is a partial order if it is also anti-symmetric. If, in addition, any two elements are related, then it is a total order. $\preceq$ is called the resource ordering. A paradigmatic example is that of entanglement ordering of bipartite quantum states, where the free channels consist of local operations and classical communication [25]. In this sense, a resource theory (and its associated resource ordering) is defined by a choice of resource objects and of free operations [26]. For instance, if the resource objects studied are quantum states, then the set of free operations is typically a subset of quantum channels closed under composition.
In our case, resource objects are instead physical theories, since (non)contextuality is a property of theories rather than states. However, as contextuality can be witnessed by individual systems within the theory, we restrict our investigation to individual systems. That is, we identify each resource object as a system in a general probabilistic theory (GPT) [27, 28, 29, 30], such as classical probability theory, quantum theory, or a subtheory thereof like the stabilizer subtheory [31]. The GPT system specifies all possible probabilistic behaviours of this physical system and thus characterizes its information-theoretic properties.
Since our resources are not states of a physical system, but rather physical theories, this implies that the transformations that we consider cannot be standard physical operations, as is the case for traditional resource theories of quantum states. Our operations of interest are simulations, which are transformations between theories that faithfully encode the information of one theory within another. Simulations that preserve indistinguishability have been introduced in [32] as univalent simulations. Since noncontextuality of an ontological model means that the model corresponds to a univalent simulation, one can show [32] that a GPT system is noncontextual if and only if it can be simulated in a univalent way by a classical GPT system — one, whose states form a simplex of probability distributions on a finite sample space. Classical GPT systems thus cannot generate contextuality under univalent simulations. Therefore, we propose the following hierarchy of (generalized) contextuality:
A GPT system $\mathsf{B}$ is said to be at least as contextual as a GPT system $\mathsf{A}$ is if there exists a univalent simulation of $\mathsf{A}$ by a composite of $\mathsf{B}$ and a classical GPT system.
It is the resource ordering of a resource theory with objects given by GPT systems and free operations given by univalent simulations with free access to classical systems. In Section 5.1, we discuss a possible interpretation of (non-univalent) simulations as expressing a particular kind of information erasure.
Contextuality monotones.
As is common in resource theories, the hierarchy of contextuality is not a total order — there are GPT systems such that neither is at least as contextual as the other one. It is not even a partial order, because there exist distinct GPT systems which are equivalent. For example, this is the case for all noncontextual systems, such as the classical bit and the classical trit. Therefore, the hierarchy is given by a preorder and it cannot be fully represented by a single numerical value — the “amount of contextuality” — assigned to each GPT system. However, in order to capture certain aspects of the hierarchy it is useful to define quantities which are order-preserving assignments of a number to each resource object. These are called resource monotones.
We define a new contextuality monotone that we call classical excess. It expresses the minimal error of a univalent simulation of a given GPT system by any classical system. In addition, we show that the optimal success probability for the parity-oblivious-multiplexing (POM) protocol [6] with free classical systems is a monotone.
Previous works on the matter.
Resource-theoretic perspective on contextuality has been developed in several works in the past. Most of these focus on the Kochen–Specker notion of contextuality [33]. Despite being related to the generalized notion of contextuality (Kochen–Specker noncontextuality is the conjunction of measurement noncontextuality and outcome determinism for sharp measurements), Kochen–Specker contextuality favors certain frameworks that, instead, are not appropriate for developing a resource theory of generalized noncontextuality. The work of Abramsky, Barbosa, and Mansfield [34] uses a framework whose main objects are empirical models — tables of data, specifying probability distributions over the joint outcomes of sets of compatible measurements. This framework is further developed in [35] and [36] and is based on the sheaf theoretic approach to contextuality introduced in [37]. Existing quantifications of Kochen–Specker contextuality are based on the memory cost [38], the ratio of contextual assignments [39], the relative entropy and contextual cost [40], the contextual robustness [41], the contextual fraction [34], and the rank of contextuality [42]. In [43] a review of several of the previous approaches towards a resource theory of Kochen–Specker contextuality is presented.
The first work on a resource theory of generalized contextuality in prepare-and-measure scenarios was presented by Duarte and Amaral in [44]. They use the generalized-noncontextual polytope characterizing the contextual set of prepared-and-measured statistics defined in [45] to motivate the set of free operations and then define monotones based on known resource quantifiers for contextuality and nonlocality. As an application of such a resource theory, [46] uses it to simplify and robustify proofs of contextuality.
On the use of general probabilistic theories.
Unlike Duarte and Amaral, we use the framework of GPTs. A GPT system consists of a collection of states, effects and a probability assigned to each pair of a state and an effect. The so-called simplex embeddability condition [47] states that a GPT system is noncontextual if and only if the associated GPT system can be embedded within a classical GPT system. This makes the GPT framework a suitable one to study generalized contextuality. The simplex embeddability condition thus characterizes the qualitative divide between contextual and noncontextual GPT systems. Our hierarchy of contextuality can be seen as a refinement thereof. It makes distinctions between contextual systems based on how (and how much) they violate simplex embeddability. In practice, this is achieved employing the notion of simulations of GPT systems from other GPT systems, as defined in [32]. We dedicate Section 5.2 to the discussion of the relation with other works that study contextuality in GPTs, namely [32, 44, 47, 48, 49, 50, 51, 52].
Structure of the paper.
In Section 2, we recall the standard treatment of generalized noncontextuality in the framework of operational theories and ontological models. This is connected to the language of general probabilistic theories in Section 3. There, we also describe (univalent) simulations of GPT systems and the excess measure. The hierarchy of contextuality is introduced in Section 4, where we prove that the classical excess is a resource monotone (Theorem 26) and construct a monotone from the parity-oblivious-multiplexing protocol. In Section 5, we discuss a possible interpretation of the non-free operations as involving information erasure and the relation to previous works on contextuality and GPTs. We conclude with a summary of the results and an outline of possible future research directions in Section 6.
2 Operational theories, ontological models and contextuality
In this work we consider prepare-and-measure scenarios associated with a single system. An operational theory associated with a prepare-and-measure scenario is defined by a list of possible preparations, measurements and the probabilities $\mathcal{P}(k|P,M)$ of obtaining the outcome $k$ of the measurement $M$ given that the system is prepared in the preparation $P$. An ontological model of an operational theory provides a realist explanation of the operational predictions of the theory [53]. It does so by stipulating the existence of an ontic state space for each given system, denoted with $\Lambda$, which is mathematically represented by a (standard Borel) measurable set. Each point $\lambda\in\Lambda$ represents an ontic state that describes all the physical properties of the system. The ontological model associates each preparation $P$ in the operational theory with a conditional probability distribution $\mu_{P}(\lambda)\equiv\mu(\lambda|P)$ over ontic states. We call these epistemic states as they represent states of knowledge about the underlying ontic states. Each outcome $k$ given a measurement $M$, denoted by $[k|M]$, is associated with a conditional probability distribution $\xi(k|\lambda,M)$. The latter corresponds to the probability of obtaining outcome $k$ given that measurement $M$ is implemented on a system in the ontic state $\lambda$. An ontological model of an operational theory reproduces the predictions of the theory via the classical law of total probability,
$\mathcal{P}(k|P,M)$ $=\sum_{\lambda\in\Lambda}\xi(k|\lambda,M)\mu(\lambda|P).$ (1)
An ontological model is preparation noncontextual if operationally equivalent preparation procedures are represented by identical probability distributions in the ontological model [2]. More formally, two preparation procedures $P$ and $P^{\prime}$ are operationally equivalent if they provide the same operational statistics for all possible measurements, i.e., $\forall M:\mathcal{P}(k|P,M)=\mathcal{P}(k|P^{\prime},M)$. In this case, we write $P\simeq P^{\prime}$. An ontological model is preparation noncontextual if any two such preparations are represented by the same epistemic states:
$P\simeq P^{\prime}\implies\mu_{P}=\mu_{P^{\prime}}.$ (2)
Similarly, two measurement outcomes $[k|M]$ and $[k^{\prime}|M^{\prime}]$ are operationally equivalent if they give the same statistics for all possible preparations: $\mathcal{P}(k|M,P)=\mathcal{P}(k^{\prime}|M^{\prime},P)$ for all preparations $P$. In this case, we write $[k|M]\simeq[k^{\prime}|M^{\prime}]$. An ontological model is measurement noncontextual if any two such measurement outcomes are represented by the same response functions:
$[k|M]\simeq[k^{\prime}|M^{\prime}]\implies\xi_{[k|M]}=\xi_{[k^{\prime}|M^{\prime}]}.$ (3)
An operational theory is termed preparation noncontextual (resp. measurement noncontextual) if there exists a preparation noncontextual (resp. measurement noncontextual) ontological model for the theory, while it is termed preparation contextual (resp. measurement contextual) if it does not admit of a preparation noncontextual (resp. measurement noncontextual) ontological model. It was first proven in [2] that quantum theory is preparation contextual and measurement noncontextual when outcome determinism is not assumed.
We note that the notion of generalised contextuality extends to transformations also [2]. In the present work we will be concerned only with prepare-and-measure scenarios and as such we will ignore transformation contextuality.
3 Contextuality for general probabilistic theories
One can often prove contextuality of a theory by studying statistical behaviours of an individual system $\mathsf{A}$ (such as a qubit) in prepare-and-measure scenarios. The information that suffices to describe the possible behaviours constitutes of
  • a set $\Omega$ of possible states that $\mathsf{A}$ can be prepared in, according to the theory in question,
  • a set of possible measurements that can be applied to $\mathsf{A}$, each consisting of a collection of effects associated with the measurement outcomes, and
  • for every state $\omega\in\Omega$ and every effect $e$ the probability of obtaining $e$ when measurement $M$ is applied to system $\mathsf{A}$ prepared in state $\omega$.
This information is commonly expressed as a system in a general probabilistic theory — a GPT system.
In the following we omit any mention of dynamics that the system may undergo. This simplification has no consequence on our discussion of preparation and measurement contextuality.
3.1 GPT systems
Let us now discuss the mathematical description of GPT systems [54, 55, 56, 57] and the definition of contextuality in this context.
Definition 1.
A GPT system $\mathsf{A}$ is specified by two non-empty convex sets
$\Omega_{\mathsf{A}}\subseteq V_{\mathsf{A}},\qquad E_{\mathsf{A}}\subseteq V_{\mathsf{A}}^{*},$ (4)
where $V_{\mathsf{A}}$ is a normed vector space and $V_{\mathsf{A}}^{*}$ is the topological dual of $V_{\mathsf{A}}$. The spaces $\Omega_{\mathsf{A}}$ and $E_{\mathsf{A}}$ are called the state space and effect space, respectively, while their elements are states and effects of $\mathsf{A}$. We require that $E_{\mathsf{A}}$ contains the null effect $0\in E_{\mathsf{A}}$ and the unit effect $1\in E_{\mathsf{A}}$, which are the 0 and 1 constant functionals, respectively, in $V_{\mathsf{A}}^{*}$.
[…]
$e\cdot\omega\in[0,1],$ (5)
for all effects $e$ and all states $\omega$. Finally, to avoid complications arising from degenerate systems, we require that
$V_{\mathsf{A}}=\mathrm{span}(\Omega_{\mathsf{A}})\qquad\text{and}\qquad V_{\mathsf{A}}^{*}=\mathrm{span}(E_{\mathsf{A}})$ (6)
hold, so that states can be distinguished by effects:
$e\cdot\omega=e\cdot\omega^{\prime}\quad\forall\,e\in E_{A}\qquad\implies\qquad\omega=\omega^{\prime},$ (7)
and similarly that effects can be distinguished by states.
The set of all GPT systems is denoted by $\mathfrak{G}$ and the subset of finite-dimensional ones by $\mathfrak{G}_{\mathrm{fin}}$.
A GPT system encodes all the information about the statistical behaviour of a physical system in prepare-and-measure scenarios. Let us give a few examples of GPT systems.
Example 2 (Finite-dimensional quantum system as a GPT system).
To each Hilbert space $\mathbb{C}^{n}$ with $n\in\mathbb{N}$, one can associate a quantum GPT system $\mathsf{Q}_{n}$. The real vector space $V_{\mathsf{Q}_{n}}\simeq\mathbb{R}^{n^{2}}$ is the space of Hermitian operators on the underlying Hilbert space. The state space is the set of density operators and the effect space is the set of quantum effects, which are positive semi-definite operators $e$ that satisfy
$0\leq e\leq 1,$ (8)
where $1$ is the identity operator. The pairing is nothing but the Hilbert–Schmidt inner product:
[…]
Likewise, every classical probabilistic system is a GPT system [54, §4].
Example 3 (Finite classical GPT systems).
We denote the finite $n$-level classical system (as well as its state space) by $\Delta_{n}$, with underlying vector space $\mathbb{R}^{n}$ and state space given by the simplex
[…]
where $\{\delta_{i}\}_{i}$ is a chosen orthonormal basis of $\mathbb{R}^{n}$ and $\mathrm{Conv}$ denotes the convex hull operation. The effect space consists of all the linear functionals $\xi$ that satisfy
$\xi\cdot\mu\in[0,1]$ (11)
for every distribution $\mu\in\Delta_{n}$. If we think of $\mu$ as a column vector, then the possible effects are all the associated row vectors with entries in the unit interval $[0,1]$. In other words, the effect space is
[…]
where
[…]
and $\{\delta^{*}_{i}\}_{i}$ is the dual basis to the one above.
Besides finite classical systems, we will also need the notion of a countably infinite one.
Example 4 (Countable classical GPT system).
The countable classical system $\Delta_{\mathbb{N}}$ has associated vector space $\ell^{1}$, the Banach space of sequences whose series are absolutely convergent. The state space, also denoted by $\Delta_{\mathbb{N}}$, consists of all probability measures on $\mathbb{N}$. The effect space $\Delta_{\mathbb{N}}^{*}$ consists of arbitrary sequences in $[0,1]^{\mathbb{N}}$ and forms a convex subset of $\ell^{\infty}$, the space of bounded sequences, which is the dual of $\ell^{1}$.
Besides individual systems, we will occasionally also need to refer to composite ones. In general, the composite of two GPT systems is not unique – see [58, Section 5] for an in-depth discussion of tensor products of GPT systems. However, among all the possible choices, there is a ‘minimal composite’ of two GPT systems $\mathsf{A}$ and $\mathsf{B}$, whose state and effect space merely contain the separable ones. Any meaningful composite of two GPT systems necessarily contains their minimal composite as a subsystem.
Definition 5.
Given two GPT systems $\mathsf{A}=(\Omega_{\mathsf{A}},E_{\mathsf{A}},V_{\mathsf{A}})$ and $\mathsf{B}=(\Omega_{\mathsf{B}},E_{\mathsf{B}},V_{\mathsf{B}})$ we define the minimal composite system $\mathsf{A}\otimes\mathsf{B}$ to be given by the separable states and separable effects on $\mathsf{A}$ and $\mathsf{B}$. Formally $\mathsf{A}\otimes\mathsf{B}:=(\Omega_{\mathsf{A}\otimes\mathsf{B}},E_{\mathsf{A}\otimes\mathsf{B}},V_{\mathsf{A}\otimes\mathsf{B}})$ where,
$V_{\mathsf{A}\otimes\mathsf{B}}$ $\coloneqq V_{\mathsf{A}}\otimes V_{\mathsf{B}},$ (14)
[…]
Of particular interest is the tensor product of an arbitrary GPT system and a classical GPT system, in which case there is a unique choice of a sensible tensor product corresponding to the minimal composite of GPT systems.
3.2 Operational theory associated to a GPT system
As we articulate in Section 2, proofs of (preparation) contextuality involve distinct preparation procedures, which are nevertheless operationally equivalent. Since any two distinct states of a GPT system are operationally inequivalent, we cannot identify GPT states with preparation procedures. Instead, we think of a preparation procedure as an ensemble of GPT states $\omega_{i}$, each occurring with a specified probability $\lambda_{i}$.
Definition 6.
A preparation $P$ of a GPT system is a finite ensemble
$\bigl{(}\lambda_{i},\omega_{i}\bigr{)}_{i},$ (17)
where $(\lambda_{i})_{i}$ is a probability distribution. It has the same statistical behaviour as the GPT state
$\omega_{P}=\sum_{i}\lambda_{i}\omega_{i},$ (18)
which is called the mean of the preparation. Two preparations, $P=(\lambda_{i},\omega_{i})_{i}$ and $P^{\prime}=(\lambda^{\prime}_{i},\omega^{\prime}_{i})_{i}$, are said to be equivalent, denoted $P\sim P^{\prime}$, if their means are equal as GPT states. We denote the set of all preparations of a GPT system $\mathsf{A}$ by $\mathrm{Prep}(\mathsf{A})$.
Similarly, proofs of measurement contextuality require the notion of distinct measurement outcomes which are operationally equivalent. Let us now give the definition of measurements and measurement outcomes for GPT systems.
Definition 7.
A measurement $M$ of a GPT system $\mathsf{A}$ is a finite collection $\{e_{1},\ldots,e_{n}\}$ of effects that is normalized, i.e. it satisfies
$\sum_{k=1}^{n}e_{k}=1.$ (19)
The set of all finite ensembles of measurements of $\mathsf{A}$ is denoted by $\mathrm{Meas}(\mathsf{A})$, which has elements $(\lambda_{i},M_{i})_{i}$ where $(\lambda_{i})_{i}$ is a probability distribution.
Given such a measurement, the probability of observing outcome $k$ provided that system $\mathsf{A}$ is prepared in state $\omega$ is given by
$\mathrm{Pr}(k|M,\omega)=e_{k}\cdot\omega.$ (20)
Definition 8.
A measurement outcome $[k|M]$ of a GPT system is a pair of an effect $e_{k}$ and a measurement $M=\{e_{1},\ldots,e_{n}\}$ which contains $e_{k}$. To stress that the effect $e_{k}$ is thought of as part of the measurement $M$, we denote it by $e_{k|M}$. Two measurement outcomes $[k|M]$ and $[k^{\prime}|M^{\prime}]$ are equivalent, denoted $[k|M]\sim[k^{\prime}|M^{\prime}]$, whenever they are equal as GPT effects:
$e_{k|M}=e_{k^{\prime}|M^{\prime}}.$ (21)
The set of all ensembles of measurement outcomes of a GPT system $\mathsf{A}$ has elements $(\lambda_{i},e_{i})_{i}$, with $(\lambda)_{i}$ a probability distribution, and is denoted by $\mathrm{Out}(\mathsf{A})$. [^4] An ensemble of outcomes $\bigl{\{}\bigl{(}\lambda,[k|M]\bigr{)},\bigl{(}1-\lambda,[k^{\prime}|M^{\prime}]\bigr{)}\bigr{\}}$ is constructed by considering the ensemble of measurements $\{(\lambda,M),(1-\lambda,M^{\prime})\}$ and identifying the outcomes $[k|M]$ and $[k^{\prime}|M^{\prime}]$. The effect corresponding to this new outcome is $\lambda e_{k|M}+(1-\lambda)e_{k^{\prime}|M^{\prime}}$. The operational equivalences of the set can be defined in analogy to Definition 6.
Remark 9 (GPT system defined by an operational theory).
State and effect spaces defining a GPT system can be constructed from an operational theory specified by a collection of preparation and measurement procedures [59], which are list of instructions to be carried out in the lab, as well as a set of all possible conditional probabilities of a measurement outcome given a preparation. In this approach, states are defined to be equivalence classes of preparation procedures, where two preparation procedures are equivalent if they cannot be distinguished using any available measurement procedure. Similarly effects are defined as equivalence classes of measurement outcomes. While this construction is more general (for instance it might include preparation procedures with extraneous information), the only preparations considered in proofs of preparation contextuality are those that correspond to preparations as introduced in Definition 6 and similarly for proofs of measurement contextuality and measurement outcomes as defined in Definition 8.
One is often particularly interested in preparations supported on the extremal points of the convex set of states $\Omega$. For classical systems, extremal states are point distributions. For quantum systems, they are the rank 1 density operators. While in the classical case there are no non-trivial equivalences among preparations of extremal states, in quantum theory they do arise.
3.3 Simulations of GPT systems
Recall from Section 2 that a physical system is noncontextual if it admits of a noncontextual ontological model. An ontological model is traditionally defined for an operational theory, from which the GPT system can be derived, as we discussed in Remark 9. In this article, we associate a canonical operational theory to every GPT system, by endowing it with the space $\mathrm{Prep}(\mathsf{A})$ of preparation procedures, $\mathrm{Meas}(\mathsf{A})$ of measurement procedures and $\mathrm{Out}(\mathsf{A})$ of measurement outcomes.
This allows us to define an ontological model for a GPT system.
Definition 10.
Given a GPT system $\mathsf{A}=(\Omega_{\mathsf{A}},E_{\mathsf{A}},V_{\mathsf{A}})$ a discrete ontological model of $\mathsf{A}$ is given by a finite classical system $\Delta_{n}$ and maps
$\begin{split}\mu:\mathrm{Prep}(\mathsf{A})&\to\Delta_{n},\\ \mu:P&\mapsto\mu_{\mathcal{P}},\end{split}\qquad\qquad\begin{split}\chi:\mathrm{Out}(\mathsf{A})&\to\Delta_{n}^{*},\\ \chi:[k|M]&\mapsto\chi_{k|M_{j}}\end{split}$ (22)
satisfying
$\chi_{k|M}\cdot\mu_{P}=e_{k|M}\cdot\omega_{P}$ (23)
for any preparation $P\in\mathrm{Prep}(\mathsf{A})$ and any measurement outcome $[k|M]\in\mathrm{Out}(\mathsf{A})$, where $\omega_{P}$ is the mean of $P$. The ontological model is called preparation noncontextual if for any pair of equivalent preparations $P$ and $P^{\prime}$, we have
$\mu_{P}=\mu_{P^{\prime}}.$ (24)
Similarly the ontological model is called measurement noncontextual if for any pair of equaivalent outcomes $[k|M]$ and $[k^{\prime}|M^{\prime}]$ we have
$\chi_{k|M}=\chi_{k^{\prime}|M^{\prime}}.$ (25)
A model which is both measurement and preparation noncontextual is a noncontextual model.
Definition 11.
If $\mathsf{A}$ admits of a discrete noncontextual ontological model, we say that it is a noncontextual GPT system. We denote the collection of all such systems by $\mathfrak{G}_{\mathtt{NC}}$. A system for which all ontological models are contextual is called a contextual system.
A noncontextual ontological model represents the behaviour of the chosen system by the behaviour of a classical system, while preserving operational equivalences (Definitions 6 and 8). It is, in fact, a special case of the following notion of simulation [32].
Definition 12.
Given two GPT systems $\mathsf{A}$ and $\mathsf{B}$, an $\epsilon$-simulation of $\mathsf{A}$ by $\mathsf{B}$ is a pair of
  • a multi-valued function $\Gamma:\Omega_{\mathsf{A}}\to\Omega_{\mathsf{B}}$ called state simulation and
  • a multi-valued function $\Theta:E_{\mathsf{A}}\to E_{\mathsf{B}}$ called effect simulation,
which is empirically adequate up to an error $\epsilon\in[0,\infty)$:
$\big{\lvert}e\cdot\omega-f\cdot\gamma\big{\rvert}\leq\epsilon\qquad\forall\,f\in\Theta(e),\;\forall\,\gamma\in\Gamma(\omega),$ (26)
and which preserves convex mixtures: [^5] Analogously, we require that the same holds for mixtures of effects rather than states, just as in [32].
$\lambda\Gamma(\omega)+(1-\lambda)\Gamma(\omega^{\prime})\subseteq\Gamma\bigl{(}\lambda\omega+(1-\lambda)\omega^{\prime}\bigr{)}\qquad\lambda\in[0,1]\text{ and }\omega,\omega^{\prime}\in\Omega_{\mathsf{A}},$ (27)
and the 00 effect:
$0_{\mathsf{B}}\in\Theta(0_{\mathsf{A}}).$ (28)
A 00-simulation is called an exact simulation.
Proposition 13 (Theorem 1 from [32]).
A discrete ontological model of a GPT system $\mathsf{A}$ provides an exact simulation of $\mathsf{A}$ by a finite classical GPT system (and vice versa).
We note that continuous ontological models of GPT systems correspond to simulations by infinite dimensional systems. However, for the purposes of this work and in keeping with the related approaches of [47, 32], we restrict ourselves to discrete ontological models.
Given a preparation $P=(\lambda_{i},\omega_{i})_{i}$ of the GPT system $\mathsf{A}$ and a simulation as above, preparations of $\mathsf{B}$ consistent with $P$ are the elements of the following subset of $\Omega_{\mathsf{B}}$:
[…]
Given an outcome $[k|M]=(\lambda_{i},e_{i})_{i}$ of the GPT system $\mathsf{A}$ and a simulation as above, outcomes of $\mathsf{B}$ consistent with $[k|M]$ are the elements of the following subset of $E_{\mathsf{B}}$:
[…]
We can use this notion to generalize the noncontextuality property of ontological models to the case of simulations following [32].
Definition 14.
A simulation of GPT systems is termed preparation univalent if for any two equivalent preparations $P$ and $P^{\prime}$, we have
$Q\sim Q^{\prime}$ (31)
for all $Q\in\Gamma(P)$ and all $Q^{\prime}\in\Gamma(P^{\prime})$.
A simulation of GPT systems is termed measurement univalent if for any two equivalent measurement outcomes $[k|M]$ and $k^{\prime}|M^{\prime}$, we have
$[q|W]\sim[q^{\prime}|W^{\prime}]$ (32)
for all $[q|W]\in\Theta([k|M])$ and all $[q^{\prime}|W^{\prime}]\in\Theta([k^{\prime}|M^{\prime}])$.
A simulation which is both preparation and measurement univalent is a univalent simulation.
Example 15 (Simulations of a bit by a trit).
Given a bit $\Delta_{2}$ and a trit $\Delta_{3}$ one can define a univalent simulation $(\Gamma,\Theta)$ of $\Delta_{2}$ by $\Delta_{3}$ via a direct embedding. Namely, the state and effect simulations act on the bases of $\mathbb{R}^{2}$ and its dual (see Example 3) via
$\Gamma\left(\delta_{i}\right)=\left\{\delta_{i}\right\},\qquad\Theta\left(\delta_{i}^{*}\right)=\left\{\delta_{i}^{*}\right\}$ (33)
and can be extended to all states and effects by linearity. This simulation is univalent since both $\Gamma$ and $\Theta$ are single-valued, a connection shown below in Proposition 18.
One can also adapt it to a simulation that is not preparation univalent while still being measurement univalent. For instance, we can keep the action on $\delta_{1}$, $\delta_{1}^{*}$ and adjust the simulation of the other deterministic state and effect to be
[…]
Once we require that $\Theta$ also preserves the unit effect, we can extend both to $\Delta_{2}$ and $\Delta_{2}^{*}$ by convex-linearity (and using Minkowski sum for the state simulation).
Example 16 (Holevo-Beltrametti-Bugasjki model [60, 59]).
Given a GPT system $\mathsf{A}=(\Omega_{\mathsf{A}},E_{\mathsf{A}},V_{\mathsf{A}})$ with $n\in\mathbb{N}$ extremal states, the Holevo-Beltrametti-Bugasjki model is given by the following simulation $(\Gamma,\Theta):\mathsf{A}\to\Delta_{n}$ where
$\Gamma\left(\omega_{i}\right)=\left\{\delta_{i}\right\},\qquad\Theta\left(e\right)=\left\{\sum_{i=1}^{n}(e\cdot\omega_{i})\delta_{i}^{*}\right\}$ (35)
where $\omega_{i}$ is the $i$-th extremal state of $\Omega_{\mathsf{A}}$ and $e$ is an arbitrary effect in $E_{\mathsf{A}}$. However, it is not a univalent simulation unless $\Omega_{\mathsf{A}}$ is a simplex. While the effect simulation is clearly single-valued, to extend the state simulation to all states of $\Omega_{\mathsf{A}}$ we need to ensure that property (27) holds. That is, for a given state $\omega$, its simulation $\Gamma(\omega)$ is given by the set of all convex decompositions of $\omega$ into extremal states (and applying $\Gamma$ to each extremal state to land within $\Delta_{n}$).
In [32] Müller and Garner define a univalent simulation to be single-valued. Our opinion is that Definition 14, stipulating preservation of operational equivalence, is more directly connected to generalized contextuality. Nevertheless, the two definitions are equivalent by Proposition 18 below. As a result, univalent $\epsilon$-simulations can be identified with the notion of $\epsilon$-embeddings [32, Definition 2], for which both the state and effect simulations are (single-valued) linear maps. Indeed, one can show that every univalent $\epsilon$-simulation extends to an essentially unique $\epsilon$-embedding and vice versa [32, Lemma 2].
Definition 17.
Given a pair of GPT systems $(\mathsf{A},\mathsf{B})$, we say that $\mathsf{A}$ is embeddable within $\mathsf{B}$, denoted by $\mathsf{A}\hookrightarrow\mathsf{B}$, if there exists an exact univalent simulation (equivalently a 00-embedding) of type $\mathsf{A}\to\mathsf{B}$. The relation $\hookrightarrow$ is called the embeddability preorder.
Indeed, one can easily show that $\hookrightarrow$ is a preorder relation. Specifically, the identity maps provide embeddings of type $\mathsf{A}\hookrightarrow\mathsf{A}$, which proves its reflexivity. To prove that it is a transitive relation, one can use the triangle inequality (see Proposition 21).
Proposition 18 (Theorem 1 [32]).
A simulation is preparation univalent if and only if its state simulation is single-valued. Similarly, a simulation is measurement univalent if and only if its effect simulation is single valued.
Proof.
First, assume that $\Gamma$ is single-valued. Then, for every preparation $P$, the set $\Gamma(P)$ has exactly one element, which is
$\bigl{(}\lambda_{i},\Gamma(\omega_{i})\bigr{)}.$ (36)
The simulation thus preserves preparation means by Equation 27, and is therefore preparation univalent.
On the other hand, consider any simulation that is not single-valued. That is, there is a state $\omega\in\Omega_{\mathsf{A}}$ such that there are two distinct states $\sigma_{1},\sigma_{2}\in\Omega_{\mathsf{B}}$ satisfying
$\{\sigma_{1},\sigma_{2}\}\subseteq\Gamma(\omega).$ (37)
Applying the state simulation to the preparation given by the delta distribution on $\omega$ gives
$\{\delta_{\sigma_{1}},\delta_{\sigma_{2}}\}\subseteq\Gamma(\delta_{\omega}).$ (38)
However, the means of the two preparations on the left are distinct — they are given by $\sigma_{1}$ and $\sigma_{2}$ respectively. Thus, the simulation is not preparation univalent.
An analogous proof holds for measurement univalence and single-valuedness of the effect simulation. ∎
3.4 Properties of univalent simulations
The following proposition, which stresses how the composition with a classical system does not affect the univalency of a simulation, is relevant for motivating our notion of contextuality preorder further on (Definition 23) as well as for quantifying contextuality.
Proposition 19.
Let $n$ and $k$ be arbitrary natural numbers. If there is a univalent $\epsilon$-simulation $(\Gamma,\Theta):\mathsf{A}\otimes\Delta_{n}\to\mathsf{B}$, then there is also a univalent $\epsilon$-simulation $\mathsf{A}\to\mathsf{B}\otimes\Delta_{k}$.
Proof.
We show this for $\epsilon=0$, the general case is analogous. In particular, we define maps $\Gamma_{1}:\Omega_{\mathsf{A}}\to\Omega_{\mathsf{B}}\otimes\Delta_{k}$ and $\Theta_{1}:E_{\mathsf{A}}\to E_{\mathsf{B}}\otimes\Delta_{k}^{*}$ via
$\Gamma_{1}(\omega)\coloneqq\Gamma(\omega\otimes\delta_{1})\otimes\delta_{1},\qquad\qquad\Theta_{1}(e)\coloneqq\Theta(e\otimes\delta^{*}_{1})\otimes\delta^{*}_{1}.$ (39)
It follows that this is an exact univalent simulation by the assumption that $(\Gamma,\Theta)$ is. Specifically, we have
$\Theta(e\otimes\delta^{*}_{j})\cdot\Gamma(\omega\otimes\delta_{i})=(e\cdot\omega)(\delta^{*}_{j}\cdot\delta_{i})=\begin{dcases}e\cdot\omega&\text{if }i=j\\ 0&\text{otherwise,}\end{dcases}$ (40)
which implies $\Theta_{1}(e)\cdot\Gamma_{1}(\omega)=e\cdot\omega$. ∎
By determining the smallest error $\epsilon$ for which there is a univalent $\epsilon$-simulation of type $\mathsf{A}\to\mathsf{B}$, we obtain a meaningful notion of how far $\mathsf{A}$ is from being embeddable within $\mathsf{B}$.
Definition 20.
The excess is the function $\varepsilon:\mathfrak{G}\times\mathfrak{G}\to\mathbb{R}$ defined by
[…]
We call $\varepsilon(A,B)$ the $A$-excess within $B$.
Note that the value of excess is always in the interval $[0,1]$. One can show that it cannot exceed $1$ by observing that there is always a univalent $1$-simulation of type $\mathsf{A}\to\Delta_{1}$ and a univalent 00-simulation of type $\Delta_{1}\to\mathsf{B}$, and using the following proposition.
Proposition 21.
For any three GPT systems $A$, $B$, and $C$, we have the triangle inequality for excess
$\varepsilon(A,C)\leq\varepsilon(A,B)+\varepsilon(B,C).$ (42)
Proof.
Consider an arbitrary univalent $\epsilon$-simulation $(\Gamma,\Theta):A\to B$ and an arbitrary univalent $\gamma$-simulation $(\alpha,\beta):B\to C$. We prove the statement by showing that $(\alpha\,\Gamma,\beta\,\Theta)$ is a univalent simulation of type $A\to C$ with error $\epsilon+\gamma$, a result stated as Lemma 3 in [32].
To show that the composite simulation is empirically adequate as expressed by Inequality (26), we can use the triangle inequality. In particular, for all $e_{\mathsf{A}}\in E_{A}$ and all $\omega_{\mathsf{A}}\in\Omega_{\mathsf{A}}$, we have
$\big{\lvert}e_{\mathsf{A}}\cdot\omega_{\mathsf{A}}$ $-\beta\,\Theta(e_{\mathsf{A}})\cdot\alpha\,\Gamma(\omega_{\mathsf{A}})\big{\rvert}\leq$ (43)
$\leq\big{\lvert}e_{\mathsf{A}}\cdot\omega_{\mathsf{A}}-\Theta(e_{\mathsf{A}})\cdot\Gamma(\omega_{\mathsf{A}})\big{\rvert}+\big{\lvert}\Theta(e_{\mathsf{A}})\cdot\Gamma(\omega_{\mathsf{A}})-\beta\,\Theta(e_{\mathsf{A}})\cdot\alpha\,\Gamma(\omega_{\mathsf{A}})\big{\rvert}$
$\leq\epsilon+\gamma,$
where the second inequality follows because $\Theta(e_{\mathsf{A}})$ is an effect of $B$ and $\Gamma(\omega_{\mathsf{A}})$ is a state of $B$, by assumption. ∎
Note that whenever $\mathsf{B}$ is embeddable within $\mathsf{C}$, we have $\varepsilon(\mathsf{B},\mathsf{C})=0$ and the triangle inequality reads
$\varepsilon(\mathsf{A},\mathsf{C})\leq\varepsilon(\mathsf{A},\mathsf{B})$ (44)
for any GPT system $\mathsf{A}$. A similar statement follows by setting the first term on the right-hand side of inequality (42) to zero. These two statements can be summarized as the following corollary of Proposition 21.
Corollary 22.
Let $\mathsf{A}$ and $\mathsf{C}$ be arbitrary GPT systems. With respect to the embeddability preorder among GPT systems, we have the following order-preserving functions:
[…]
$\varepsilon(\mathsf{A},\mathsf{C})>\varepsilon(\mathsf{B},\mathsf{C})$ (47)
holds (i.e. the minimal error of simulation by $C$ is strictly higher for $\mathsf{A}$ than for $\mathsf{B}$), then there cannot be an exact univalent simulation of type $\mathsf{A}\to\mathsf{B}$. We can thus use the excess within $\mathsf{C}$, for any GPT system $\mathsf{C}$, as a witness of the impossibility to construct an exact univalent simulation.
While simulations are crucial to express a binary division between noncontextual and contextual GPT systems, the embeddability preorder $\hookrightarrow$ is not a satisfactory refinement of this binary division. The simplest argument to see this is that all classical GPT systems (which are trivially noncontextual) should be equivalent in a resource theory of contextuality. However, for any integer $m$ greater than $k$, the classical GPT system $\Delta_{m}$ is not embeddable within $\Delta_{k}$. Therefore, we cannot think of the embeddability preorder $\hookrightarrow$ as a resource ordering of some resource theory of contextuality. In the following section, we motivate a preorder that further refines the embeddability relation $\hookrightarrow$ and addresses the issue above by stipulating classical GPT systems to be ‘free resources’.
4 Hierarchy of contextuality
Recall that our interest in a hierarchy of contextuality of GPT systems stems from the wish to measure how useful contextuality is as a resource in information-theoretic applications. How can one decide whether a given measure of interest is indeed a well-motivated measure of contextuality? The resource-theoretic perspective [61, 26], fruitfully applied to studies of entanglement and many other resources in the past, suggests to first motivate a preorder $\succeq$ among the objects of interest. Importantly, some objects may be incomparable according to $\succeq$. In our case, the resource objects are GPT systems and the resource ordering is the advertised hierarchy of contextuality. The interpretation of such an ordering is that the relation $\mathsf{A}\succeq\mathsf{B}$ expresses the statement ‘$\mathsf{A}$ is at least as contextual as $\mathsf{B}$ is’. Given a good motivation for such a hierarchy, one can answer the above question as follows:
A (real-valued) function on the set $\mathfrak{G}$ of all GPT systems is a measure of contextuality if and only if it is order-preserving with respect to the hierarchy of contextuality $\succeq$.
4.1 Motivation and the resource theory
Let us now list and discuss our desiderata for a sensible hierarchy of contextuality among GPT systems. They are:
  1. 1.
    Access to classical GPT systems is free. That is, we have
    $\mathsf{A}\preceq\mathsf{A}\otimes\Delta_{n}\quad\text{ and }\quad\mathsf{A}\otimes\Delta_{n}\preceq\mathsf{A}$ (48)
    for every GPT system $\mathsf{A}$ and every classical GPT system $\Delta_{n}$. Here, $\otimes$ denotes the minimal composite from Definition 5.
  2. 2.
    Exact univalent simulations are free. That is, if $\mathsf{A}$ is embeddable within $\mathsf{B}$, then we have $\mathsf{A}\preceq\mathsf{B}$.
Notice that the first condition in Equation 48, i.e. $\mathsf{A}\preceq\mathsf{A}\otimes\Delta_{n}$, follows from desideratum 2 because every GPT system $\mathsf{A}$ is embeddable within $\mathsf{A}\otimes\Delta_{n}$ (see Proposition 19).
There are several consequences of desiderata 1 and 2 which we want to point the reader’s attention to, before we discuss our motivation to introduce these desiderata.
  1. (i)
    Desideratum 2 implies that the trivial GPT system $\Delta_{1}$ is least contextual in the sense that it is the bottom element of $\preceq$. That is, we have $\Delta_{1}\preceq\mathsf{A}$ for every GPT system $\mathsf{A}$.
  2. (ii)
    Consequently, using the fact that $\Delta_{1}\otimes\Delta_{n}$ is the same GPT system as $\Delta_{n}$, every classical GPT system is also at the bottom by Property (1).
  3. (iii)
    Using definition 11 and desideratum 2, we thus get that all noncontextual GPT systems are equivalent to each other and lie at the bottom of the hierarchy.
Motivation for Requirement 1.
Contextuality, as we are thinking of it, is a property that expresses the degree to which a system’s behaviour escapes a purely classical explanation. In this sense, considering additional classical systems should not affect this property. An important point to note is that within the tensor product $\mathsf{A}\otimes\Delta_{n}$ (Definition 5), only correlations involving separable states between the generic GPT system $\mathsf{A}$ and the classical system are allowed. We can think of behaviours of $\mathsf{A}\otimes\Delta_{n}$ as dilations of those of $\mathsf{A}$ [62], given a classical environment variable $\Delta_{n}$.
If there were classical environmental variables, whose knowledge reduces the contextuality present in $\mathsf{A}$, then the contextuality of $\mathsf{A}$ would not be an authentic feature of $\mathsf{A}$, rather a consequence of one’s too narrow focus on $\mathsf{A}$ as opposed to its dilation $\mathsf{A}\otimes\Delta_{n}$. We can also support this argument on the technical side. Indeed, by Proposition 19, any noncontextual model of $\mathsf{A}\otimes\Delta_{n}$ provides also a noncontextual model of $\mathsf{A}$. In other words, $\mathsf{A}$ is noncontextual if and only if $\mathsf{A}\otimes\Delta_{n}$ is. [^6] Note that under the minimal tensor product of Definition 5 allowing access to arbitrary noncontextual systems (and not just classical systems) would lead to the same hierarchy, which follows from the fact that access to classical systems already places all noncontextual systems at the bottom of the hierarchy. The hierarchy of contextuality ought to capture this fact.
Motivation for Requirement 2.
We can also think of contextuality (of a GPT system) as the inability to provide a noncontextual model (Definition 11) for its statistical behaviours. If a GPT system $\mathsf{A}$ can be exactly simulated by a GPT system $\mathsf{B}$ via a univalent simulation, then any obstruction to constructing such a noncontextual model for $\mathsf{A}$ must already be present in $\mathsf{B}$. In this sense, we think of $\mathsf{B}$ as being “at least as contextual” as $\mathsf{A}$ is. Indeed, provided with a noncontextual model of $\mathsf{B}$, i.e. an exact univalent simulation of $\mathsf{B}$ by a classical GPT system, one can construct a noncontextual model of $\mathsf{A}$ by composition with the simulation of $\mathsf{A}$ by $\mathsf{B}$.
The preorder relation among finite-dimensional GPT systems in $\mathfrak{G}_{\mathrm{fin}}$ that satisfies precisely these requirements is the following one.
Definition 23.
We say that $\mathsf{B}$ is at least as contextual as $\mathsf{A}$, denoted $\mathsf{A}\preceq\mathsf{B}$, whenever there is an exact univalent simulation of $\mathsf{A}$ by $\mathsf{B}\otimes\Delta_{n}$ for some finite-dimensional classical GPT system $\Delta_{n}$. The preordered set $(\mathfrak{G}_{\mathrm{fin}},\succeq)$ is called the hierarchy of contextuality.
In other words, we define the hierarchy via embeddability given an additional classical system:
$\mathsf{A}\preceq\mathsf{B}\quad\iff\quad\exists\,n\in\mathbb{N}\text{ such that }\mathsf{A}\hookrightarrow\mathsf{B}\otimes\Delta_{n}\text{ holds.}$. (49)
To show that the hierarchy $\preceq$ is a preorder relation, we can use a resource-theoretic perspective, i.e. we can show that the set of free operations in the following resource theory is closed under composition.
Definition 24.
The resource theory of GPT-contextuality is defined as follows. Its objects are finite-dimensional GPT systems and each transformation is an exact simulation ${(\Gamma,\Theta):\mathsf{A}\to\mathsf{B}\otimes\Delta_{n}}$ for some $\mathsf{A},\mathsf{B}\in\mathfrak{G}_{\mathrm{fin}}$. It acts on objects by transforming $\mathsf{B}$ to $\mathsf{A}$ and is undefined for any GPT system other than $\mathsf{A}$. Transformations
$(\Gamma,\Theta):\mathsf{A}\to\mathsf{B}\otimes\Delta_{m}\quad\text{and}\quad(\alpha,\beta):\mathsf{B}\to\mathsf{C}\otimes\Delta_{k}$ (50)
can be composed (sequentially) to produce a transformation
$(\Gamma\,\alpha,\Theta\,\beta):\mathsf{A}\to\mathsf{C}\otimes\Delta_{m\cdot k}$ (51)
by using the isomorphism between $\Delta_{m}\otimes\Delta_{k}$ and $\Delta_{m\cdot k}$. Free transformations are those ones that are also univalent simulations.
This gives rise to a concrete resource theory in the sense of [26, Definition 3.10]. Moreover, the resulting resource ordering coincides with our hierarchy of contextuality. Note that in a resource theory the existence of a free operation transforming a resource object $r$ to a resource object $s$ defines the resource ordering denoted by $s\preceq r$. Thus, to match our definition of the hierarchy of contextuality, the transformation as in (50) has to be understood as a free operation from $\mathsf{B}$ to $\mathsf{A}$ (and not from $\mathsf{A}$ to $\mathsf{B}$ as the notation might suggest). This is consistent with the interpretation of a simulation of $\mathsf{A}$ by $\mathsf{B}$ whereby an experimenter has access to $\mathsf{B}$ and uses it to simulate $\mathsf{A}$, hence transforming their access to system $\mathsf{B}$ to their effective access to system $\mathsf{A}$.
4.2 Contextuality of composite systems
When studying a resource theory, it is often also important to specify how resource objects are composed to obtain a joint resource. As discussed in Section 3.1, GPT systems do not have a canonical tensor product. However, any sensible choice of a tensor product $\mathbin{\tilde{\otimes}}$ should give rise to a symmetric monoidal category with exact simulations as morphisms [58, Section 5.2]. Since the transformations in our resource theory of GPT-contextuality are not merely simulations of GPT systems, but also involve additional classical systems, it takes a bit more work to show that they (and the free transformations) are closed under parallel composition. We prove this in the rest of this section for the case of free operations.
Consider two exact univalent simulations of types $\mathsf{A}_{1}\to\mathsf{B}_{1}\otimes\Delta_{m}$ and $\mathsf{A}_{2}\to\mathsf{B}_{2}\otimes\Delta_{k}$ respectively. The assumption that $\mathbin{\tilde{\otimes}}$ is the monoidal product of a symmetric monoidal category then means that there is also an exact univalent simulation of type
$\mathsf{A}_{1}\mathbin{\tilde{\otimes}}\mathsf{A}_{2}\to\left(\mathsf{B}_{1}\otimes\Delta_{m}\right)\mathbin{\tilde{\otimes}}\left(\mathsf{B}_{2}\otimes\Delta_{k}\right).$ (52)
Assuming that there is a unique composite of any GPT system with a classical one given by the canonical composite of Definition 5 (see [63] for an alternative composition of classical systems) and that $\mathbin{\tilde{\otimes}}$ is associative, the codomain can be expressed as
$\begin{split}\left(\mathsf{B}_{1}\otimes\Delta_{m}\right)\mathbin{\tilde{\otimes}}\left(\mathsf{B}_{2}\otimes\Delta_{k}\right)&=\left(\mathsf{B}_{1}\mathbin{\tilde{\otimes}}\Delta_{m}\right)\mathbin{\tilde{\otimes}}\left(\mathsf{B}_{2}\mathbin{\tilde{\otimes}}\Delta_{k}\right)\\ &\cong\left(\mathsf{B}_{1}\mathbin{\tilde{\otimes}}\mathsf{B}_{2}\right)\mathbin{\tilde{\otimes}}\left(\Delta_{m}\mathbin{\tilde{\otimes}}\Delta_{k}\right)\\ &\cong\left(\mathsf{B}_{1}\mathbin{\tilde{\otimes}}\mathsf{B}_{2}\right)\mathbin{\tilde{\otimes}}\Delta_{m\cdot k}\\ &=\left(\mathsf{B}_{1}\mathbin{\tilde{\otimes}}\mathsf{B}_{2}\right)\otimes\Delta_{m\cdot k},\end{split}$ (53)
where the third equality uses the standard isomorphism between $\Delta_{m}\otimes\Delta_{k}$ and $\Delta_{m\cdot k}$. In the first and final steps we also use the assumption that composition with a classical system is unique. Consequently, we get a parallel composite of the two free transformations as an exact univalent simulation of type
$\mathsf{A}_{1}\mathbin{\tilde{\otimes}}\mathsf{A}_{2}\to\left(\mathsf{B}_{1}\mathbin{\tilde{\otimes}}\mathsf{B}_{2}\right)\otimes\Delta_{m\cdot k}.$ (54)
As a consequence, the parallel composition of resources (i.e. GPT systems) via $\mathbin{\tilde{\otimes}}$ respects the hierarchy of contextuality in the sense that we have
$\mathsf{A}_{1}\preceq\mathsf{B}_{1}\;\land\;\mathsf{A}_{2}\preceq\mathsf{B}_{2}\quad\implies\quad\mathsf{A}_{1}\mathbin{\tilde{\otimes}}\mathsf{A}_{2}\preceq\mathsf{B}_{1}\mathbin{\tilde{\otimes}}\mathsf{B}_{2},$ (55)
for arbitrary GPT systems $\mathsf{A}_{i}$ and $\mathsf{B}_{i}$. Condition (55) is part of the definition of a resource theory via ordered commutative monoids in [64]. It means that we can consistently apply the hierarchy of contextuality also to express the contextuality of composite GPT systems, as long as their composition satisfies a few basic properties as described above.
4.3 Quantifying contextuality via the classical excess
As is common in resource theories [65], we can use monotones — i.e. order-preserving functions from resources to numbers — to study the hierarchy of contextuality. On the one hand, these can provide a lens through which to extract properties of the preorder. On the other hand, they can be used as quantitative measures of contextuality. The latter perspective is particularly useful when the monotones come equipped with an (operational) interpretation that allows one to identify which aspect of contextuality they are measuring.
In order to construct a specific contextuality monotone, we will make use of the excess function from Definition 20. Of particular interest is $\varepsilon(\mathsf{A},\Delta_{\mathbb{N}})$, the $\mathsf{A}$-excess within the countable classical GPT system (Example 4). For brevity, we call it the classical excess of $\mathsf{A}$. A crucial ingredient in the following proof is that $\Delta_{\mathbb{N}}$ is isomorphic to $\Delta_{\mathbb{N}}\otimes\Delta_{n}$ for any finite $n$.
Lemma 25.
For any GPT system $\mathsf{A}$ and any finite-dimensional classical GPT system $\Delta_{n}$, we have
$\varepsilon(\mathsf{A}\otimes\Delta_{n},\Delta_{\mathbb{N}})=\varepsilon(\mathsf{A},\Delta_{\mathbb{N}})$ (56)
Proof.
Firstly, the inequality
$\varepsilon(\mathsf{A}\otimes\Delta_{n},\Delta_{\mathbb{N}})\geq\varepsilon(\mathsf{A},\Delta_{\mathbb{N}})$ (57)
is a direct consequence of Proposition 19 and the definition of excess.
For the converse inequality, note that any $\epsilon$-embedding $(\phi,\psi):\mathsf{A}\to\Delta_{\mathbb{N}}$ can be extended to an $\epsilon$-embedding
$(\phi\otimes\mathrm{id},\psi\otimes\mathrm{id})\;:\;\mathsf{A}\otimes\Delta_{n}\to\Delta_{\mathbb{N}}\otimes\Delta_{n},$ (58)
so that we get
$\varepsilon(\mathsf{A},\Delta_{\mathbb{N}})\geq\varepsilon(\mathsf{A}\otimes\Delta_{n},\Delta_{\mathbb{N}}\otimes\Delta_{n}).$ (59)
Furthermore, there is a 00-embedding $\Delta_{\mathbb{N}}\otimes\Delta_{n}\to\Delta_{\mathbb{N}}$ given by the linear map
$\ell^{1}\otimes\mathbb{R}^{n}$ $\to\ell^{1}$ (60)
$(\omega_{i})_{i\in\mathbb{N}}\otimes\delta_{j}$ $\mapsto\bigl{(}\omega_{i/j}\delta_{j|i}\bigr{)}_{i\in\mathbb{N}},$
where $(\omega_{i})_{i\in\mathbb{N}}$ is an element of $\ell^{1}$, $\delta_{j}$ is a basis vector in $\mathbb{R}^{n}$, and $\delta_{j|i}$ returns $1$ if $j$ divides $i$ and 00 otherwise. Using basically the same linear map for effects, merely extended to $\ell^{\infty}\otimes\mathbb{R}^{n}\to\ell^{\infty}$, gives the required embedding. Consequently, we have
$\varepsilon(\Delta_{\mathbb{N}}\otimes\Delta_{n},\Delta_{\mathbb{N}})=0,$ (61)
so that by inequality (59) and Proposition 21 we get
$\varepsilon(\mathsf{A},\Delta_{\mathbb{N}})$ $\geq\varepsilon(\mathsf{A}\otimes\Delta_{n},\Delta_{\mathbb{N}}\otimes\Delta_{n})+\varepsilon(\Delta_{\mathbb{N}}\otimes\Delta_{n},\Delta_{\mathbb{N}})$ (62)
$\geq\varepsilon(\mathsf{A}\otimes\Delta_{n},\Delta_{\mathbb{N}})$
as required. ∎
Theorem 26.
[…]
$\mathsf{A}\preceq\mathsf{B}\quad\implies\quad\varepsilon(\mathsf{A},\Delta_{\mathbb{N}})\leq\varepsilon(\mathsf{B},\Delta_{\mathbb{N}}).$ (63)
Proof.
By Definition 23, Corollary 22 (specialized to $\mathsf{C}=\Delta_{\mathbb{N}}$), and Lemma 25 respectively, we have
$\mathsf{A}\preceq\mathsf{B}$ $\iff\quad\exists\,n\;\text{ such that }\;\mathsf{A}\hookrightarrow\mathsf{B}\otimes\Delta_{n},$ (64)
$\;\implies\quad\exists\,n\;\text{ such that }\;\varepsilon(\mathsf{A},\Delta_{\mathbb{N}})\leq\varepsilon(\mathsf{B}\otimes\Delta_{n},\Delta_{\mathbb{N}})$
$\iff\quad\varepsilon(\mathsf{A},\Delta_{\mathbb{N}})\leq\varepsilon(\mathsf{B},\Delta_{\mathbb{N}}).$
This gives us the statement of the theorem. ∎
Remark 27 (Excess within a finite simplex).
An important feature of the monotone $\varepsilon(\mathsf{A},\Delta_{\mathbb{N}})$ is that it can sometimes be computed by considering the $\mathsf{A}$-excess within $\Delta_{k}$ for a finite $k$. This happens whenever its value $\varepsilon(\mathsf{A},\Delta_{m})$ remains constant for all classical systems with $m$ larger than $k$. For instance, the noncontextual GPT system corresponding to the toy bit in Spekkens toy theory [66] is embedabble within $\Delta_{4}$ with vanishing error, and hence for any larger $m$ its excess within $\Delta_{m}$ has to remain 00.
4.4 Parity oblivious multiplexing success probability with free classical resources as a measure of contextuality
Having introduced the hierarchy of contextuality as well as a new contextuality measure in the form of the classical excess, we now build a measure of generalised contextuality based on the parity oblivious multiplexing (POM) [10] game. The latter is a protocol that is powered by preparation contextuality and has raised significant attention in recent years [67, 68, 69, 70, 71, 72, 73, 74].
In the $\bm{n}$-bit Parity Oblivious Multiplexing (POM) game, Alice is given an $n$-bit string, whose possible values she encodes in a GPT system $\mathsf{A}$. In particular she chooses a state $\omega_{x}$ for possible string value $x$. The system is then transmitted to Bob is additionally given an integer $y\in\{1,2,\ldots,n\}$. Bob’s task is to guess the value $x_{y}$ of the yth bit of $x$ by performing a two-outcome measurement $M_{y}$. His guess is denoted by $b\in\{0,1\}$.
The task has a supplementary constraint called parity obliviousness (PO). Namely, Alice cannot communicate the parity of the string, denoted by $p(x)$, to Bob. The PO condition can be phrased in terms of Alice’s chosen states as
$\sum_{x|p(x)=0}\omega_{x}=\sum_{x|p(x)=1}\omega_{x}.$ (65)
Note that the string ${x}$ is assumed to be uniformly distributed, so that the left-hand side of Equation 65 can be interpreted (up to a factor of $2^{n-1}$) as the average state sent by Alice given that the characters of ${x}$ add up to 00 modulo $2$ and similarly for the right-hand side.
The average success probability is given by
$\sum_{x,y}\frac{1}{2^{n}n}p(b=x_{y}\,|\,\omega_{x},M_{y}),$ (66)
where $2^{-n}$ is the probability of Alice receiving string $x$, $1/n$ is the probability of Bob receiving $y$, and ${p(b=x_{y}\,|\,\omega_{x},M_{y})}$ is the probability that Bob’s guess is correct in a given run.
It is shown in [10, Theorem 2] that if a system $\mathsf{A}$ is noncontextual, then the optimal (i.e. maximal) success probability for the $n$-bit POM, denoted by $\mathfrak{p}_{n}$, is upper bounded by
$\mathfrak{p}_{n}(\mathsf{A})\leq\frac{n+1}{2n}=\mathfrak{p}_{n}(\Delta_{\infty}).$ (67)
The success probability for POM can be thus used as a witness of contextuality.
We now prove that for the embeddability preorder the POM success probability is also order-preserving.
Lemma 28.
Let $n$ be any integer. If $\mathsf{A}$ is embeddable within $\mathsf{B}$, then we have $\mathfrak{p}_{n}(\mathsf{A})\leq\mathfrak{p}_{n}(\mathsf{B})$.
Proof.
Let us consider any choice of states $\omega_{x}$ (for $x\in\{0,1\}^{n}$) which obey the parity obliviousness constraint (65). Since we have $\mathsf{A}\hookrightarrow\mathsf{B}$, for any choice of effects $e_{y}$ and states $\omega_{x}$ on $\mathsf{A}$ obeying the parity obliviousness constraint, the states $\Gamma(\omega_{x})$ on $\mathsf{B}$ also obey it by the linearity of the single-valued state simulation $\Gamma$. Moreover, the empirical adequacy (Equation 26) of the exact simulation implies $e_{y}\cdot\omega_{x}=\Gamma(e_{y})\cdot\Gamma(\omega_{x})$. Thus, any strategy on $\mathsf{A}$ can be implemented on $\mathsf{B}$ with equal success probability. This implies $\mathfrak{p}_{n}(\mathsf{A})\leq\mathfrak{p}_{n}(\mathsf{B})$. ∎
However, it is not a resource monotone in our resource theory of GPT-contextuality.
Lemma 29.
The optimal success probability for POM is not a monotone with respect to the hierarchy of contextuality $\preceq$.
Proof.
We define a noisy bit to be any system with the same states as the classical bit $\Delta_{2}$, but whose effects form a proper subset of $\Delta_{2}^{*}$. In particular, a noisy bit has no measurement that distinguishes the two extremal states $\delta_{1}$ and $\delta_{2}$ with certainty. Noisy bits can be parametrized by a noise parameter $\alpha\in(0,1/2]$, giving rise to a GPT system $\Delta_{2}^{\alpha}$ whose extremal effects $\chi_{1}$ and $\chi_{2}$ satisfy
$\chi_{i}\cdot\delta_{j}=\begin{cases}1-\alpha&\text{if }i=j\\ \alpha&\text{if }i\neq j.\end{cases}$ (68)
By the desiderata 1 and 2, a noisy bit $\Delta_{2}^{\alpha}$ and a bit $\Delta_{2}$ must satisfy $\Delta_{2}^{\alpha}\preceq\Delta_{2}$ and $\Delta_{2}\preceq\Delta_{2}^{\alpha}$ since they are both noncontextual. The optimal strategy for POM given a (noisy) bit consists of encoding the first bit $x_{1}$ of the string $x$ in its extremal states $\delta_{1}$ and $\delta_{2}$ [10, Lemma 1]. With probability $1/n$ Bob is asked to guess this specific bit, which he can do successfully with probability $(1-\alpha)$. With probability $(n-1)/{n}$ Bob is asked to guess any of the remaining bits from $x$, for which he can do no better than a uniformly random guess, i.e. the success probability is $1/2$. Hence, the optimal success probability is
$\mathfrak{p}_{n}(\Delta_{2}^{\alpha})=\frac{1-\alpha}{n}+\frac{n-1}{2n}=\frac{n+1-2\alpha}{2n},$ (69)
which is strictly smaller than $\mathfrak{p}_{n}(\Delta_{2})=(n+1)/2n$ whenever the noise parameter $\alpha$ is greater than 00. This means that $\Delta_{2}\preceq\Delta_{2}^{\alpha}$ does not imply $\mathfrak{p}_{n}(\Delta_{2}^{\alpha})\geq\mathfrak{p}_{n}(\Delta_{2})$. ∎
We can see that the POM success probability cannot serve as a monotone because the POM game does not allow the use of classical systems for free. This can be remedied by the so-called generalized yield construction [65, Section 3.1].
Definition 30.
Given a GPT system $\mathsf{A}$ we define the optimal POM success probability with free classical resources as:
$\mathfrak{p}_{n}{\textrm{-}}\mathrm{yield}(\mathsf{A})\coloneqq\sup_{d\in\mathbb{N}}\mathfrak{p}_{n}(\mathsf{A}\otimes\Delta_{d}).$ (70)
Proposition 31.
$\mathfrak{p}_{n}{\textrm{-}}\mathrm{yield}$ is a monotone on the resource theory of GPT-contextuality.
Proof.
Let us assume $\mathsf{A}\preceq\mathsf{B}$, which implies that there is an $m\in\mathbb{N}$ such that there is a 00-embedding $\mathsf{A}\otimes\Delta_{d}\to\mathsf{B}\otimes\Delta_{d\cdot m}$ for every $d\in\mathbb{N}$. By Lemma 28, we obtain $\mathfrak{p}_{n}(\mathsf{A}\otimes\Delta_{d})\leq\mathfrak{p}_{n}(\mathsf{B}\otimes\Delta_{d\cdot m})$ for every $d\in\mathbb{N}$, which yields the inequality below:
$\sup_{d\in\mathbb{N}}\mathfrak{p}_{n}(\mathsf{A}\otimes\Delta_{d})\leq\sup_{d\in\mathbb{N}}\mathfrak{p}_{n}(\mathsf{B}\otimes\Delta_{d\cdot m}).$ (71)
Moreover,
$\sup_{d\in\mathbb{N}}\mathfrak{p}_{n}(\mathsf{B}\otimes\Delta_{d\cdot m})=\sup_{d\in\mathbb{N}}\mathfrak{p}_{n}(\mathsf{B}\otimes\Delta_{d}),$ (72)
which implies
$\mathfrak{p}_{n}{\textrm{-}}\mathrm{yield}(\mathsf{A})\leq\mathfrak{p}_{n}{\textrm{-}}\mathrm{yield}(\mathsf{B}),$ (73)
using Definition 30. ∎
This proof is an instance of the proof of the generalized yield construction in [26, Theorem 4.21], so that $\mathfrak{p}_{n}{\textrm{-}}\mathrm{yield}$ can be meaningfully interpreted as a (generalized) yield monotone.
5 Discussion
5.1 Contextuality and information erasure
The fine-tuning problem of contextuality.
Contextuality of a theory implies the existence of distinctions at the ontological level which are not present at the operational level. If a contextual ontological model truly describes the physical reality underlying the observed behaviours predicted by the theory, then there are operationally indistinguishable behaviours that have distinct ontological origins. In other words, such operational equivalences would result from a fine-tuning of the corresponding distinct ontological representations [4]. How do these distinctions disappear between the ontological and operational descriptions of the physical system, though? The presence of such fine-tunings provides a conspiratorial connotation to the realist explanation of the theory and we believe that it requires an explanation. In this section, we explore the possibility that the fine-tuning associated with contextuality can be explained as emergent from a yet undiscovered physical mechanism that supplements the description provided by the ontological model.
Explaining fine-tunings as emergent from yet undiscovered physical mechanisms.
Explaining the origin of fine-tunings of this kind by searching for new physical mechanisms dates back to Valentini’s veriant of Bohmian mechanics [75]. There, he introduces a notion of quantum equilibrium as the reason why superluminal signaling does not manifest in quantum theory, despite the nonlocality of its underlying ontological model. This picture predicts that outside of the quantum equilibrium, it is possible to observe faster than light signaling. Therefore, the fine-tuned nature of no-signaling in Bohmian mechanics is explained just as an emergent feature of the quantum equilibrium and it is not universally valid. We cannot avoid noticing how radical such explanations of fine-tunings rooted in undiscovered physical mechanisms are. They imply that an established physical principle, such as the principle of no-signaling, is violated at the fundamental level. In the case of contextuality, the physical mechanism explaining the emergence of the operational equivalences would entail the existence of measurements that can distinguish behaviours that are deemed indistinguishable by quantum theory.
Explaining contextuality through information erasure.
In Valentini’s work the quantum equilibration process is responsible for the emergence of no-signalling — the fine-tuned feature associated with nonlocality. What hypothetical physical mechanism could be responsible for the emergence of operational equivalences — the fine-tuned feature associated with contextuality?
It would have to be a process that involves a kind of information erasure. The information erased is the information about distinctions at the ontological (e.g. fundamental) level, which cannot be stored in systems of the operational (e.g. effective) theory that lacks these distinctions. By Landauer’s principle, we can then associate an increase in entropy between the fundamental and effective levels. Such a process of information erasure would not only provide an explanation for the problematic fine-tuning associated with contextuality but would also be associated to a potentially detectable heat dissipation. This heat would signify that, indeed, there are distinctions at the fundamental level which are not present at the effective level. One could even hypothesise that the information erasure is a physical process occuring over time. That is, during the preparation of a quantum system there may be a timescale before which the system is described by the fundamental (and noncontextual) theory. At longer timescales, once the erasure has occured, the system can only be described by the effective (contextual) theory.
Ontological model as a fundamental theory.
The above account treats an ontological model as specifying a yet unfalsified fundamental theory, in accordance with [32]. This departs from the standard use of ontological models to study contextuality [2]. In the latter, the ontological distinctions (which are not present in the operational theory) are in principle indistinguishable. On the other hand, the fundamental theory we posit here contains no such requirement. In particular, the in-principle-indistinguishability of its distinctions would also mean that there is no entropy increase and no heat resulting from erasure to be detected. Examples of effective theories arising from more fundamental ones include thermodynamics, which emerges from statistical mechanics via coarse-graining, and classical information theory, which emerges from quantum information theory via decoherence. The difference between the two interpretations of ontological models does not prevent us from using the approach of [32] to provide an explanation of generalized contextuality as defined in [2]. If the fine-tuning associated with generalized contextuality is explained through a process information erasure, then the problematic aspect of contextuality in quantum theory disappears. Instead, one is then led to search for an in principle accessible more fundamental theory, from which quantum theory emerges.
Resource theory of GPT-contextuality and physical erasure processes.
In our work, both the effective and fundamental theory are associated with GPT systems. Consider $\mathsf{A}$ as the effective description of a physical system and $\mathsf{B}$ the fundamental one. A simulation $\mathsf{A}\to\mathsf{B}$ can be then interpreted as specifying, for each behaviour of the effective theory, all behaviours of the fundamental theory that would give rise to it. For example, if the effective description arises from the fundamental one due to a physical erasure process given by some many-to-one function $f:\mathsf{B}\to\mathsf{A}$, then the associated simulation would be given by preimages under $f$. This point of view is consonant with the use of GPT simulations in [32] where they were introduced. However, not every simulation can be associated with a physical map (in the opposite direction) in such a way. Here, a physical map $\mathsf{B}\to\mathsf{A}$ would be a linear map $f:V_{\mathsf{B}}\to V_{\mathsf{A}}$ satisfying $f(\Omega_{\mathsf{B}})\subseteq\Omega_{\mathsf{A}}$, $f^{*}(E_{\mathsf{A}})\subseteq E_{\mathsf{B}}$ and $f^{*}(e_{\mathsf{A}})\cdot\omega_{\mathsf{B}}=e_{\mathsf{A}}\cdot f(\omega_{\mathsf{B}})$. For instance, while the Spekkens toy bit [66] has a noncontextual ontological model (and thus a univalent simulation) defined on the tetrahedron $\Delta_{4}$, there is no physical map from $\Delta_{4}$ to the toy bit. Some ontological models do correspond to physical maps, though. For any GPT system $\mathsf{A}$ with $n$ extremal states $\omega_{i}$ there exists a Holevo-Beltrametti-Bugasjki (HBB) model of $\mathsf{A}$ by $\Delta_{n}$, as shown in Example 16. The HBB simulation corresponds to a linear map $M:\mathbb{R}^{n}\to V_{\mathsf{A}}$ whose action on the extremal states of $\Delta_{n}$ is given by: $\delta_{i}\mapsto\omega_{i}$. See Appendix A for a proof that $M$ is a physical map.
Erasure associated with ontological models.
By analyzing physical maps like the one of the HBB simulation we can understand why information erasure in the context of ontological models and contextuality cannot be the same as the standard thermodynamic one. While the latter can be understood as a coarse-graining operation, the physical maps associated to contextual ontological models are not of this kind. To see this, consider a contextual GPT system $\mathsf{A}$ and a classical system $\Delta_{d}$ such that a physical map $f:\Delta_{d}\to\mathsf{A}$ exists. Since $\mathsf{A}$ is contextual, by Lemma 11 of [32], there is no coarse-graining (i.e. idempotent) process on $\Delta_{d}$ whose image is $\mathsf{A}$. Hence, the physical GPT map $f$ cannot be understood simply as a coarse-graining operation, such as the erasure that occurs between fundamental and effective levels in statistical mechanics. A more concrete perspective is offered by the HBB map $M$ which explicitly models the preparation of the system $\mathsf{A}$ via the classical control $\Delta_{n}$. The classical degree of freedom is used to select the extremal state of $\mathsf{A}$ to be prepared. Thus, while $M$ can be used to model a physical process in the laboratory (e.g. the preparation of $\mathsf{A}$ conditioned on a classical variable), it does not correspond to a simple coarse-graining process applied to the classical degree of freedom.
5.2 Relation with previous works on contextuality and GPTs
On the framework of Müller and Garner.
On the technical level, our work makes use of the tools developed in [32]. For instance, the concept of univalent simulations of GPT systems plays an especially important role. Even though we use similar tools, our goal is quite different. Müller and Garner use the concept of univalent simulations to formulate a notion of generalized noncontextuality that applies to generic effective and corresponding fundamental theories. The effective theory is the theory emerging from the operational statistics of the experiment and the fundamental theory is a fine-grained theory of the effective ones (so, not just the simplicial theory, as in the case of standard noncontextuality).
They use such noncontextuality as a plausibility criterion for testing the validity of the fundamental theory given the effective theory. This approach becomes particularly meaningful when the fundamental theory is assumed to be quantum theory, since it allows the authors to provide experimental tests of quantum theory. Related works which employ an alternative definition of noncontextuality are the ones of Gitton and Woods [48] and [49]. For ongoing debate on the merits of the different approaches see [49] and [76].
On simplex embeddability.
Other influential sources of inspiration for our work are [47, 51, 52]. As we already mentioned in the introduction, [47] introduces the notion of simplex embeddability as the geometrical criterion to assess whether a GPT system is noncontextual. [^7] Another work that is closely related to [47] is that of Shahandeh [50]. However, the criterion of classicality introduced there additionally requires the simplex to have the same dimension of the GPT system. We do not adopt this approach. For example, it deems the rebit stabilizer theory to be nonclassical. However, the rebit stabilizer theory, in the prepare-and-measure scenario, admits a noncontextual ontological model given by the Spekkens toy theory [66], as has been shown in several works [77, 78, 79]. The criterion is extended to accessible GPT fragments in [51], where the latter correspond to more general mathematical objects than GPTs and characterize generic prepare-and-measure experimental setups. The work of [52] provides an algorithm for testing contextuality in any prepare-and-measure scenario. In particular, if it exists, it returns an explicit noncontextual model for the scenario and, if not, it provides the minimum amount of (depolarizing) noise which would be required until a noncontextual model would become possible. They call this measure of contextuality the robustness of nonclassicality. The latter is related to our notion of error of univalent simulation, but it is different insofar as it requires a specific noise channel (e.g. depolarizing noise channel), whereas the error of univalent simulation does not require that. Moreover, the robustness of nonclassicality is not defined as a measure within a well defined resource theory.
On accessible GPT fragments.
It has been argued in [51] that in realistic experiments GPT systems are generally not appropriate to study contextuality. In particular, the following discrepancy arises: A GPT system obtained from an operational description of the experiment, as in Remark 9, provides states and effects observable in practice in this experiment, but does not limit the in principle implementable ones. However, the preparation equivalences preserved by a preparation-noncontextual ontological model (see Section 2 and Definition 10) are in principle operational equivalences.
When studying preparation contextuality only, this discrepancy can be dealt with. Namely, one can first set-up the effect space $E_{\mathsf{A}}$ in a non-operational way, as a hypothesized set of in principle possible effects. Importantly, one has to identify each actual measurement outcome arising in the experiment with a (distribution over) the elements of $E_{\mathsf{A}}$. Their means specify a subset $\tilde{E}_{\mathsf{A}}\subseteq E_{\mathsf{A}}$ of the so-called accessible effects, which however play no role in the study of preparation contextuality. The state space $\Omega_{\mathsf{A}}$ is the set of equivalence classes of preparation procedures as in Remark 9, but only with respect to in principle operational equivalence. That is, states that are indistinguishable by accessible effects, but could be distinguished by a hypothesized effect in $E_{\mathsf{A}}\setminus\tilde{E}_{\mathsf{A}}$ ought not be identified in $\Omega_{\mathsf{A}}$. With this set-up, one can study preparation contextuality as we discuss shortly in Section 3.3 in terms of the GPT system $(\Omega_{\mathsf{A}},E_{\mathsf{A}})$.
Such a treatment is not possible when studying both preparation and measurement contextuality simultaneously. One has to replace the notion of a GPT system by the notion of an accessible fragment of a GPT system introduced in [51]. Simply put, this is a pair of a GPT system $(\Omega_{\mathsf{A}},E_{\mathsf{A}})$ together with a subsystem $(\tilde{\Omega}_{\mathsf{A}},\tilde{E}_{\mathsf{A}})$ thereof, which importantly can violate the non-degeneracy condition (7), and therefore are not necessarily GPT system. The enveloping system describes states and effects that can arise in principle, while the subsystem describe those that are achievable in the particular experiment, whose contextuality we are interested in.
While it would be arguably more appropriate to develop our hierarchy of contextuality for accessible fragments instead of plain GPT systems, we chose against this option for two reasons. Firstly, this would make the approach superficially more complicated, even though all the key ideas can be presented in the simpler framework of GPT systems. Secondly, one can easily accommodate our approach to the case of accessible fragments. The only difference is that the preparation (and outcome) equivalences have to be adjusted. For example, two preparations (i.e. distributions over $\tilde{\Omega}_{\mathsf{A}}$) would be deemed equivalent if their means cannot be distinguished by any effect in $E_{\mathsf{A}}$.
On the resource theory of contextuality due to Duarte and Amaral.
Another related work is [44], where the authors also provide a resource theory of generalized contextuality in prepare-and-measure scenarios using a ‘black box’ framework. In [44] the basic object consists of a prepare-and-measure scenario $\mathcal{S}$ together with a behaviour $B$. A scenario is a set $\mathcal{S}=\{\mathcal{P},\mathcal{M},\mathcal{D},\mathcal{E}_{P},\mathcal{E}_{M}\}$ where $\mathcal{P}$ is a set of preparations, $\mathcal{M}$ a set of measurements, $\mathcal{D}$ a set of outcomes, $\mathcal{E}_{P}$ the operational equivalences for preparations and $\mathcal{E}_{M}$ the operational equivalences for measurements.
A behaviour $B$ is a set of conditional probabilities:
$B=\{p(k|j,i)\}_{i\in\mathcal{I},j\in\mathcal{J},k\in\mathcal{K}}$ (74)
corresponding to the probability of outcome $d_{k}$ given preparation $P_{i}$ and measurement $M_{j}$. $\mathcal{I},\mathcal{J}$ and $\mathcal{K}$ are sets of labels for the preparations, measurements and outcomes, respectively.
We note that the sets of operational equivalences $\mathcal{E}_{P}$ and $\mathcal{E}_{M}$ of a given object are defined independently of the probability distributions. As such an object (scenario + behaviour) may contain operational equivalences which are not implied by the behaviour. Alternatively, there may be operational equivalences present in the behaviour which are not included in the scenario. In the cases where all ensembles of procedures are valid procedures in the scenario, and where the operational equivalences implied by behaviour are those of the scenario, then the object defined is equivalent to a GPT system.
The free objects of [44] are the noncontextual behaviours, namely those which admit of a generalized noncontextual model.
The operations of [44] are given by pre- and post-processing on the preparations, measurements and outcomes and correspond to channel simulations. The free operations correspond to channel simulations such that all the operational equivalences of the simulated system are images of operational equivalences of the simulating system through the simulation map.
Let us consider an example where both resource theories can be applied to highlight the differences between the resource theory of [44] and the one presented in this work.
We consider a scenario $\mathcal{S}=\{\mathcal{P},\mathcal{M},\mathcal{D},\mathcal{E}_{P},\mathcal{E}_{M}\}$ and a scenario $\tilde{\mathcal{S}}=\{\tilde{\mathcal{P}},\tilde{\mathcal{M}},\tilde{\mathcal{D}},\tilde{\mathcal{E}}_{P},\tilde{\mathcal{E}}_{M}\}$ where,
$\mathcal{P}=\{0,1\},\quad$ =$\mathcal{M}=\{0\},\quad$ $\mathcal{D}=\{0,1\},$ $\mathcal{E}_{P}=\emptyset,$ $\mathcal{E}_{M}=\emptyset,$ (75)
$\mathcal{P}=\{0,1,2\},\quad$ =$\mathcal{M}=\{0\},\quad$ $\mathcal{D}=\{0,1,2\},\quad$ $\mathcal{E}_{P}=\emptyset,\quad$ $\mathcal{E}_{M}=\emptyset.$ (76)
The behaviours on $\mathcal{S}$ and $\tilde{\mathcal{S}}$ are $B$ and $\tilde{B}$ respectively:
$p(k|j=1,i)$ $=\delta_{i,k},\ i,k\in\{0,1\},$ (77)
$p(\tilde{k}|\tilde{j}=1,\tilde{i})$ $=\delta_{\tilde{i},\tilde{k}},\ \tilde{i},\tilde{k}\in\{0,1,2\}.$ (78)
These are the behaviours obtained by preparing a bit (resp. trit) in one of its extremal states and then implementing the canonical two outcome (resp. three outcome) measurement which returns one of the outcomes with certainty.
In the framework of GPTs $(\mathcal{S},B)$ corresponds to a bit and $(\tilde{\mathcal{S}},\tilde{B})$ to a trit (assuming one also allows all ensembles of preparations and measurements).
In the resource theory of GPT contextuality presented in this work we have that the bit and the trit are equivalent as resources: $\Delta_{2}\succeq\Delta_{3}$ and $\Delta_{3}\succeq\Delta_{2}$.
However in the resource theory of [44] this is not the case. We now show explicitly that there is no free operation $T$ taking $(\mathcal{S},B)$ to $(\tilde{\mathcal{S}},\tilde{B})$. Since there is only one measurement in both scenarios we drop the indices $j$ and $\tilde{j}$. Let us assume the existence of stochastic maps $q_{O}(\tilde{k}|k)$ and $q_{P}(i|\tilde{i})$ which map $B\to\tilde{B}$:
$p(\tilde{k}|\tilde{i})=\sum_{i,k=0}^{1}q_{O}(\tilde{k}|k)p(k|i)q_{P}(i|\tilde{i})$ (79)
Now using the definitions of $B$ and $\tilde{B}$ of Equation 77 we obtain:
$\delta_{\tilde{i},\tilde{k}}=\sum_{i=0}^{1}q_{O}(\tilde{k}|i)q_{P}(i|\tilde{i}).$ (80)
Let us denote by $Q_{O}$ the $3\times 2$ matrix with entries $[Q_{O}]_{\tilde{k},i}=q_{O}(\tilde{k}|i)$ and $Q_{P}$ the $2\times 3$ matrix with entries $[Q_{P}]_{i,\tilde{i}}=q_{P}(i|\tilde{i})$. Then the previous equation is equivalent to the matrix equation:
$\mathbb{I}_{3}=Q_{O}\cdot Q_{P}$ (81)
This yields a contradiction since the decomposition of a $3\times 3$ matrix $A$ into matrix product $A=B\cdot C$ with $B$ a $3\times 2$ matrix and $C$ a $2\times 3$ matrix implies $C$ has rank at most 2. However $\mathbb{I}_{3}$ is a rank 3 matrix. This shows that there is no free operation $(\mathcal{S},B)\mapsto(\tilde{\mathcal{S}},\tilde{B})$ in the resource theory of [44]. Notice that the reason why in our resource theory the bit and the trit — and any other noncontextual theory — are equivalent is that our free operations include the access to the classical system. This is the origin of the difference with the resource theory of [44].
An additional contribution of the present work is the following. In [44] the authors define a notion of composite of two behaviours which is equivalent to the minimal composite of Definition 5 in our framework. They then show that this composite is consistent with their resource theory, in the sense that the free objects are closed under this composite. In the present work we show that any consistent composite of GPT systems, not just the minimal composite of Definition 5 is compatible with the resource theory of contextuality.
6 Conclusion
Generalized contextuality, a leading notion of nonclassicality, is of crucial importance both in the foundations of quantum theory and in quantum information processing. Despite this, there is no complete characterisation of generalised contextuality as a resource. In this work we address this shortcoming.
Based on recent developments formulating noncontextuality of GPT systems via simplex embeddability, we have defined a resource theory of contextuality of GPT systems in prepare-and-measure scenarios. The free resources are the noncontextual systems and the free operations are univalent simulations with free access to classical systems. Using these notions, we motivate a hierarchy of contextuality for GPT systems. A new contextuality monotone arises naturally from our considerations — the classical excess, which expresses the minimum error of a univalent simulation by the countably-infinite classical system. We have also shown how a standard witness of contextuality, in the form of the POM success probability, can be used to define a contextuality monotone.
A key feature of the resource theory presented is this work is that the free operations include access to classical systems. Since every noncontextual system can be univalently simulated by a classical system this entails that all noncontextual systems are equivalent in the hierarchy. Without access to free classical systems the preorder which emerges is the embeddability preorder, for which generic noncontextual systems are inequivalent. Moreover we show, by assuming that the only way of composing GPT systems with classical ones is the standard one, that any consistent composition of GPT systems is consistent with the hierarchy: if $\mathsf{A}_{1}\preceq\mathsf{B}_{1}$ and $\mathsf{A}_{2}\preceq\mathsf{B}_{2}$ then $\mathsf{A}_{1}\mathbin{\tilde{\otimes}}\mathsf{A}_{2}\preceq\mathsf{B}_{1}\mathbin{\tilde{\otimes}}\mathsf{B}_{2}$.
We have also discussed how GPT simulations could describe a physical process of information erasure that would explain the fine-tuning associated with contextuality [^8] A recent work that connects preparation contextuality and information erasure is [80], where the authors show that any ontological model reproducing the statistics of a sequential protocol involving incompatible projective measurements involves more information erasure than what operational quantum theory predicts. This fact is strictly related to the presence of preparation contextuality, as the same final quantum state in the protocol is represented by two different ontic distributions. This result can be seen as an example of fine-tuning of information erasure (i.e. more erasure at the ontological level than at the operational level), which, despite being interesting, differs from our idea of viewing contextuality as arising from a process of information erasure that explains the operational equivalences of distinct ontological representations. in a similar way to how the quantum equilibration process proposed by Valentini explains the fine-tuning associated with nonlocality in Bohmian mechanics. We have argued that the information erasure in this case would be different than a simple coarse-graining. An interesting avenue for future research would be to further characterize this kind of information erasure and propose a test to detect heat dissipation in experiments manifesting contextuality. Even though such a proposal is undoubtedly radical, we believe that it could explain the fine-tuning associated with contextuality without abandoning the ontological models framework.
Another direction for further investigation is to develop a solid interpretation and establish the use-cases of the resource theory of GPT-contextuality that we provide. It is known that a resource theory of a given notion need not be unique. We have examples of this fact, like the resource theories of entanglement based on LOCC [81] and LOSR [82, 83] free operations, respectively. This does not mean that only one of them is “the correct resource theory of entanglement”. Rather, these resource theories may be applicable in different contexts or for different purposes. In the case of entanglement, one can say that LOCC operations are relevant for communication tasks in the context of quantum internet and LOSR operations are relevant for the study of entanglement in Bell scenarios.
In this respect, we have discussed the relation of our work with other studies of generalized contextuality for GPTs in section 5.2. One that we did not mention is [72]. It would be interesting to find a relationship between our classical excess and the simulation cost of contextuality defined therein. Finally, it would be also interesting to extend the methods developed here to the realm of resource theories of more general fine-tunings [4], such as violations of time symmetry [84] and bounded ontological distinctness [85].
Acknowledgments
The authors thank the participants of the PIMan workshop – Orange (CA), March 2019 – where the idea of this project originated. In particular, Luke Burns and Justin Dressel, who were part of the initial discussions on the project. The authors further thank Rafael Wagner for insightful explanations regarding the approach of Duarte and Amaral [44]. LC thanks Farid Shahandeh and TGa thanks Markus Müller for helpful discussions. This project started when LC was supported by the Fetzer Franklin Fund of the John E. Fetzer Memorial Trust and by the Army Research Office (ARO) (Grant No. W911NF-18-1-0178). LC also acknowledges funding from the Einstein Research Unit “Perspectives of a Quantum Digital Transformation” and from the Horizon Europe project FoQaCiA, GA no.101070558. TGa acknowledges support from the Austrian Science Fund (FWF) via project P 33730-N. TGo acknowledges support from the Austrian Science Fund. This research was funded in whole or in part by the Austrian Science Fund (FWF) via the START Prize Y1261-N. This research was supported in part by Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported by the Government of Canada through the Department of Innovation, Science, and Economic Development, and by the Province of Ontario through the Ministry of Colleges and Universities.
For open access purposes, the authors have applied a CC BY public copyright license to any accepted manuscript version arising from this submission.
Appendix A Physicality of the Holevo projection
In the framework of GPTs a physical map $\mathsf{A}\to\mathsf{B}$ is given by a linear map $M:V_{\mathsf{A}}\to V_{\mathsf{B}}$ such that $M(\Omega_{\mathsf{A}})\subset\Omega_{\mathsf{B}}$. $M$ can be viewed as part of a preparation procedure, preparing a state of $\mathsf{B}$ by first preparing a state $\omega_{\mathsf{A}}$ of $\mathsf{A}$ and then applying $M$ to obtain $\omega_{\mathsf{B}}=M\omega_{\mathsf{A}}$. Composing this with a measurement $\{e_{\mathsf{B}}^{i}\}_{i}$ on $\mathsf{B}$ gives outcome probabilities $e_{\mathsf{B}}^{i}(M\omega_{\mathsf{A}})$. However, one could also view this experiment as preparing the state $\omega_{\mathsf{A}}$ on system $\mathsf{A}$ and then measuring $\{e_{\mathsf{B}}^{i}\circ M\}_{i}=\{M^{*}e_{\mathsf{B}}^{i}\}_{i}$ which is a measurement on $\mathsf{A}$, hence we require that a physical map also obeys $M^{*}e_{\mathsf{B}}\in E_{\mathsf{A}}$ for all $e_{\mathsf{B}}\in E_{\mathsf{B}}$.
Definition 32 (Physical GPT map).
A physical map $M$ between two system $\mathsf{A}=(\Omega_{\mathsf{A}},E_{\mathsf{A}},V_{\mathsf{A}})$ and $\mathsf{B}=(\Omega_{\mathsf{B}},E_{\mathsf{B}},V_{\mathsf{B}})$ is given by a linear map $M:V_{\mathsf{A}}\to V_{\mathsf{B}}$ such that:
$M(\Omega_{\mathsf{A}})\subseteq\Omega_{\mathsf{B}},$ (82)
$M^{*}(E_{\mathsf{B}})\subseteq E_{\mathsf{A}}.$ (83)
It turns out that a (multivalent) simulation of $\mathsf{B}$ by $\mathsf{A}$ does not, in general, define a physical transformation $\mathsf{A}\to\mathsf{B}$. First, a simulation $(\Gamma,\Theta)$ from $\mathsf{B}=(\Omega_{\mathsf{B}},E_{\mathsf{B}})$ to $\mathsf{A}=(\Omega_{\mathsf{A}},E_{\mathsf{A}})$ is given by a pair of multivalued maps. The state simulation $\Gamma:\Omega_{\mathsf{A}}\to\Omega_{\mathsf{B}}$ is not, in general, a map (since it does not assign a unique element of $\Omega_{\mathsf{B}}$ to each $\Omega_{\mathsf{A}}$). The inverse $\Gamma^{-1}$: $\Gamma(\Omega_{\mathsf{A}})\subset\Omega_{\mathsf{B}}\to\Omega_{\mathsf{A}}$ defines a map $\Gamma(\Omega_{\mathsf{A}})\to\Omega_{\mathsf{A}}$, which does not, in general, linearly extend to a map $\Gamma^{-1}:V_{\mathsf{B}}\to V_{\mathsf{A}}$ and hence cannot be a physical transformation.
Moreover, in the case where $\Gamma^{-1}$ does define a linear map $V_{\mathsf{B}}\to V_{\mathsf{A}}$ it might not meet the requirement that $(\Gamma^{-1})^{*}(E_{\mathsf{A}})\subseteq E_{\mathsf{B}}$. This shows why, within the framework of GPTs, one cannot in general associate a physical process from an underlying classical system to the GPT system for a generic ontological model.
However, there is a family of simulations which do.
Lemma 33.
The HBB simulation (see Example 16) of a system $\mathsf{A}$ with finitely extremal states $\delta_{\rm ext}(\Omega_{\mathsf{A}})=X$ on the classical system $\Delta_{|X|}$ given by the multivalued maps $(\Gamma,\Theta)$ induces a physical map $\Gamma^{-1}:\Delta_{|X|}\to\mathsf{A}$.
Proof.
Given a GPT system $\mathsf{A}$ with finitely many extremal states $X$ the HBB model is given by the maps $(\Gamma,\Theta)$ as defined in Example 16. We first show that $\Gamma^{-1}:\Gamma(\Omega_{\mathsf{A}})\to\Omega(\mathsf{A})$ defines a map $\Delta_{|X|}\to\Omega_{\mathsf{A}}$.
The map $\Gamma^{-1}$ acts on extremal states of $\Delta_{|X|}$ as:
$\Gamma^{-1}:\delta_{x}\mapsto\omega_{x},x\in X.$ (84)
Extending the action to mixed states gives:
$\bar{\Gamma}^{-1}:\sum_{x}p_{x}\delta_{x}\mapsto\omega=\sum_{x}p_{x}\omega_{x}.$ (85)
This is consistent with $\Gamma:\omega\mapsto\{\sum_{x\in X}p_{x}\delta_{x}|p_{x}\in[0,1],\sum_{x\in X}p_{x}=1,\omega=\sum_{x\in X}p_{x}\omega_{x}\}$ since $\bar{\Gamma}^{-1}(\Gamma(\omega))=\omega$ for every $\mu\in\Gamma(\omega)$.
Since $\Delta_{|X|}$ spans $\mathbb{R}^{|X|}$ the map $\bar{\Gamma}^{-1}$ linearly extends to $\mathbb{R}^{|X|}$, defining a map $M:\mathbb{R}^{|X|}\to V_{\mathsf{A}}$.
This map satisfies $M^{*}(E_{\mathsf{A}})\subset\Delta_{m}^{*}$ which can be seen as follows:
$M^{*}(e_{\mathsf{A}})\omega_{\Delta_{m}}=e_{\mathsf{A}}(M\omega_{\Delta_{m}})=e_{\mathsf{A}}(\omega_{\mathsf{A}})\in[0,1].$ (86)
Hence $M^{*}(e_{\mathsf{A}})$ takes $\Delta_{m}$ to $[0,1]$ and, since $\Delta_{m}$ is unrestricted, $M^{*}(e_{\mathsf{A}})\in\Delta_{m}^{*}$. ∎